Computational Construction of W-graphs of Hecke algebras H(q, n) for n up to 15

Mitsuyuki Ochiai Fujio Kako

Abstract

We construct by computer all W-graphs corresponding to irreducible representations of Hecke algebras H(q, n) for n up to 15, using a modification of a method proposed by Lascoux and Schützenberger (which fails for n > 13).